A physical fitness association is including the mile run in its secondary-school fitness test. The time for this event for boys in secondary school is known to possess a normal distribution with a mean of 470 seconds and a standard deviation of 60 seconds. Find the probability that a randomly selected boy in secondary school will take longer than 332 seconds to run the mile.
Answer #1

Solution :

Given that ,

mean = \mu = 470

standard deviation = \sigma = 60

P(x > 332) = 1 - P(x < 332)

= 1 - P((x - \mu ) / \sigma < (332 - 470) / 60)

= 1 - P(z < -2.3)

= 1 - 0.0107   

= 0.9893

Probability = 0.9893